OpenStudy (anonymous):

on a map 3 inches represents 240 miles. five inches represents 400 miles. how many inches would be needed to represent 320 miles

OpenStudy (anonymous):

This was better. \[3i=240m \implies i=80m\] \[5i=400m \implies i=80m\] \[x*i=320m\]

OpenStudy (anonymous):

3inches=240miles 1inch=80miles It's the same for the five inches rep 400miles So , 320miles= (1/80)x320=4 Another easier way of saying, 1mile=3/240 320miles=(3/240)x320=4

OpenStudy (anonymous):

4 inches